# Efficient Redundancy Reduction for Open-Vocabulary Semantic Segmentation

### v1

@@ IV-C6 Impact of CLIP architecture
Following previous methods [15, 20], our ERR-Seg also adopts the hierarchical CLIP (ConvNeXt) instead of the plain CLIP (ViT). The impact of the CLIP architecture is illustrated in Table XII. To ensure a fair comparison, training and evaluation are conducted using an input size of $384^{2}$ with frozen CLIP. The results show that employing hierarchical CLIP leads to a +$2.2$ +and +$4.1$ mIoU improvement on A-847 and A-150, respectively. In comparison to plain CLIP, hierarchical CLIP offers the following advantages: (1) it allows the use of middle-layer features to enhance the reconstruction of high-resolution semantic maps, and (2) it permits direct inference on high-resolution images, eliminating the need for the sliding window inference strategy [16] employed in plain CLIP.
IV-C7 Impact of sampled class number
Table XIII illustrates the influence of the sampled class number $P$ in the CRM during the inference phase. It is evident that as $P$ increases gradually, the performance tends to reach a saturation point. This finding aligns with the expectation that a sufficient increase in $P$ can adequately cover all existing categories. Specifically, we assign the following values for $P$: 48 for A-847, 48 for PC-459, 32 for A-150, 24 for PC-59, and 16 for PAS-20.
IV-C8 Impact of input size
@@
Table XIV presents the results with varying input sizes during the inference phase. The findings highlight a decrease in performance when input sizes are too large or too small. Specifically, for models based on ConvNeXt-B and ConvNeXt-L, the recommended input sizes are $640^{2}$ and $768^{2}$, respectively. Our analysis reveals that using ConvNeXt-L as the backbone of CLIP, there is a better alignment between the image and text spaces, enabling the extraction of more refined pixel-text cost maps and thus supporting larger input sizes.
TABLE XIV: The impact of crop size during inference.
@@
IV-D Qualitative Results
Fig. 8 presents a qualitative comparison among SAN [14], CAT-Seg [16], and our proposed ERR-Seg on the validation sets of ADE20K-150 and ADE20K-847. The results highlight the superiority of our method over previous approaches in mask generation and open-vocabulary recognition. In addition, ERR-Seg can identify targets that were previously overlooked in ADE20K. While the current OVSS model shows remarkable open-vocabulary recognition capability due to the advancements of VLMs, it struggles with segmenting intricate masks. There is still potential for improvement in effectively recognizing small and nested targets.
@@ IV-D Qualitative Results
V Conclusion
In conclusion, we introduce ERR-Seg, a novel framework for open-vocabulary semantic segmentation (OVSS). ERR-Seg investigates the impact of redundant classes and reveals that they lead to long latency and hinder the attention mechanism from capturing spatial-level and class-level dependencies. To address this issue, ERR-Seg proposes the training-free Channel Reduction Module (CRM) to eliminate redundant classes based on prior knowledge from VLMs. Additionally, recognizing the importance of hierarchical semantics for semantic segmentation, ERR-Seg introduces the Hierarchical Semantic Module (HSM) to generate hierarchical cost maps from the open-vocabulary visual encoder. Experimental results demonstrate that ERR-Seg establishes a new state-of-the-art in terms of efficiency and accuracy. Future work will focus on exploring high-quality real-time OVSS architecture design.

### v2

@@ 4.3.5 Impact of CLIP architecture
Following previous methods [xie2024sed, yu2024convolutions], our ERR-Seg also adopts the hierarchical CLIP (ConvNeXt) instead of the plain CLIP (ViT). The impact of the CLIP architecture is illustrated in Table 13. To ensure a fair comparison, training and evaluation are conducted using an input size of $384^{2}$ with frozen CLIP. The results show that employing hierarchical CLIP leads to a +$2.2$ +and +$4.1$ mIoU improvement on A-847 and A-150, respectively. Furthermore, compared to the ViT-based architecture, it slightly reduces inference latency, with an 8 ms reduction on ADE20K-150. In comparison to plain CLIP, hierarchical CLIP offers the following advantages: (1) it allows the use of middle-layer features to enhance the reconstruction of high-resolution semantic maps, and (2) it permits direct inference on high-resolution images, eliminating the need for the sliding window inference strategy [cho2024cat] employed in plain CLIP.
Table 13: Impact of CLIP Architecture.
CLIP Arch. Latentcy A-847 PC-459 A-150 PC-59 PAS-20
ViT-B/16 56 7.5 13.0 22.8 49.7 91.0
ConvNeXt-B 47 9.7 12.5 26.9 49.4 89.9
4.3.6 Impact of input size
@@
Table 14 presents the results with varying input sizes during the inference phase. The findings highlight a decrease in performance when input sizes are too large or too small. Specifically, for models based on ConvNeXt-B and ConvNeXt-L, the recommended input sizes are $640^{2}$ and $768^{2}$, respectively. Our analysis reveals that using ConvNeXt-L as the backbone of CLIP, there is a better alignment between the image and text spaces, enabling the extraction of more refined pixel-text cost maps and thus supporting larger input sizes.
Table 14: The impact of crop size during inference.
@@
4.3.7 Failure case analysis.
Fig. 5 presents a qualitative analysis of ERR-Seg’s performance. The results indicate that our model can identify small objects, such as the person and cars depicted in Fig. 5 (a). However, the model struggles to segment precise masks for extremely narrow or small objects, such as the table legs and road railings shown in Fig. 5 (a) and (b). Moreover, ERR-Seg demonstrates considerable robustness in parsing complex scenes, such as cluttered indoor environments. Nevertheless, it remains susceptible to semantic misclassification, for instance, confusing a sofa with an armchair in Fig. 5 (a). It is worth noting that addressing these issues remains a key objective in the field of semantic segmentation.
5 Conclusion
In conclusion, we introduced ERR-Seg, an efficient framework for open-vocabulary semantic segmentation. We identified that the redundancy in cost-based methods not only increases computational latency but also hinders the attention mechanism from capturing spatial-level and class-level dependencies. To address this, ERR-Seg incorporates the Redundancy-Reduced Hierarchical Cost maps (RRHC), which eliminates irrelevant classes using semantic priors from VLMs and incorporates hierarchical cost maps for enriched semantic representation. Furthermore, the Redundancy-Reduced Cost Aggregation (RRCA) significantly reduces computational overhead via sequence compression. Experimental results demonstrate that ERR-Seg establishes a new state-of-the-art in terms of efficiency and accuracy.
Limitations and future work. Our failure case analysis reveals two primary limitations of the proposed OVSS framework: difficulties in generating precise masks for extremely small or narrow objects and inaccuracies in class recognition. We attribute these challenges to the cost aggregation paradigm’s reliance on CLIP for generating dense pixel-text cost maps, which provides insufficient fine-grained visual representations for these challenging scenarios. Future work will enhance the OVSS framework by integrating powerful vision foundation models (e.g., DINO [simeoni2025dinov3], SAM [kirillov2023segment]) to improve mask quality and leveraging advanced Multimodal Large Language Models (e.g., Qwen-VL series [Qwen-VL, Qwen2-VL, Qwen2.5-VL]) to provide more robust open-vocabulary recognition capability. Moreover, future work also will focus on designing lightweight learnable class selection module to automatically adapt to diverse vocabularies and image inputs.
Appendix A Appendix
Figure 8: Visualization of segmentation results in various domains. Our proposed ERR-Seg is capable of segmenting capybaras (a rare category in public datasets) from various domains, including (a) synthesized images, (b) cartoon images, (c) natural images, and (d) capybara dolls. Moreover, ERR-Seg achieves more precise masks than SAN [xu2023side] and CAT-Seg [cho2024cat].
A.1 Proof for Proposition 1
Proof.
$\Delta_{r}=\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}.$
Similarly, $\Delta_{p}$ can be simplified as:
$\Delta_{p}=\sum_{k=M+1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)},$
$\Delta_{r}^{\prime}=\left(\frac{1}{\Delta_{r}}-1\right),$ (14)
$\Delta_{p}^{\prime}=\left(\frac{1}{\Delta_{p}}-1\right).$ (15)
Since $0<\Delta_{r},\Delta_{p}<1$, it follows that $\Delta_{r}^{\prime},\Delta_{p}^{\prime}>0$. To compare $\Delta_{r}$ and $\Delta_{p}$, we first compare $\Delta_{r}^{\prime}$ and $\Delta_{p}^{\prime}$, by analyzing the following formula:
$\frac{\Delta_{r}^{\prime}}{\Delta_{p}^{\prime}}=\frac{\sum_{k=1}^{M}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}}{\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}}\frac{\sum_{k=M+1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}}{\sum_{k=1}^{M}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}}.$
Let us define:
$\alpha=\sum_{k=1}^{M}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{M}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}$
and
$\beta=\sum_{k=M+1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}.$
Thus, we have
$\frac{\Delta_{r}^{\prime}}{\Delta_{p}^{\prime}}=\alpha\beta.$ (19)
$\forall 1\leq i,j\leq P,A_{ij}=B_{ij}.$ (20)
This implies that $\alpha=1$ and $\beta<1$. Hence,
$\frac{\Delta_{r}^{\prime}}{\Delta_{p}^{\prime}}=\alpha\beta<1.$ (21)
@@ A.1 Proof for Proposition 1
A.2 More Qualitative Results
@@
[…]